Compute the frequency (in MHz) of an EM wave with a wavelength of 9.0 in. (0.2286 m).
1
Expert's answer
2019-05-06T10:33:39-0400

"\lambda = \frac{c}{f}"

"\lambda" - wavelength, m

"c" - speed of light, "c=3 \cdot 10^8 \; m/s"

"f" - frequency, Hz

"f = \frac{c}{\lambda}"

"f = \frac{3 \cdot 10^8}{0.2286} \approx 1312.3 \cdot 10^6 \; Hz =1312.3 \; MHz"

Answer: 1312.3 MHz.